ZZ Tire Company opened a new location last year. Use the following information to figure out contribution margin and operating income for last year:

Tires sold: 15,000

Retail Price per tire: $75 per tire

Cost per tire: $18 per tire

Installation expense per tire (payroll): $9 per tire

Utilities: $2,000 per month

Rent: $3,500 per month

Clerks and phone receptionist: $2,500 per month
Answer #1

Total Variable cost Per Unit = cost Per Unit + Installation Expenses = 18 +9 = $27

Fixed cost = Unilities + Rent clerk and Phone Receptionist = 2000+3500+2,500 = $8,000

contibution Margin = ( Selling Price -Variable cost) * Tires

Contibution Margin = (75-27) *15,000 =720,000 (Answer)

Operating income = Contibution Margin -( Fixed cost * 12)

Operating income = 720,000-(8,000*12) = $624,000 (Answer)